OpenStudy (anonymous):

what is the discriminant of 8x+2=-8x^2

OpenStudy (turingtest):

a=8 b=8 c=2 b^2-4ac plug in the numbers

OpenStudy (anonymous):

\[8x^2+8x+2=0\] \[discriminant = b^2-4(a*c)=8^2+4(8*2)=64+64=128\]

OpenStudy (turingtest):

8^2-4(8)(2)=0 you lost the "-" sign adgd

OpenStudy (turingtest):

Hear that sickdigs? The answer is 0

OpenStudy (anonymous):

oh ok thanks a lot for clarifying

OpenStudy (turingtest):

sarcasm or serious? if you need more explanation ask.

OpenStudy (anonymous):

serious i was confused but you clarified i think i got i now thanks